OpenStudy (anonymous):

Bernie counted the cars and motorcycles that passed his house in 45 minutes. He counted 12 vehicles and 36 tires. How many motorcycles passed his house? (3 points) Bernie counted the cars and motorcycles that passed his house in 45 minutes. He counted 12 vehicles and 36 tires. How many motorcycles passed his house? (3 points) @Mathematics

OpenStudy (anonymous):

Just like previous one..

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

how many cars shabeer

OpenStudy (anonymous):

explain me this one!

OpenStudy (anonymous):

x+y = 12 2x+4y = 36

OpenStudy (anonymous):

6 cars &6 motercycles

OpenStudy (anonymous):

do we have to use this one

OpenStudy (anonymous):

yes, and y = 12-x If we substitute this in second equation, we can find the value for x

OpenStudy (anonymous):

ya sheg